OpenStudy (iwanttogotostanford):

What is the equation of the following graph in vertex form? parabolic function going down from the left through the point zero comma twelve and through the point two comma zero and turning at the point four comma negative four and going up through the point six comma zero and continuing towards infinity

OpenStudy (iwanttogotostanford):

@wolf1728 @sleepyhead314

OpenStudy (sleepyhead314):

I will assume that the "turning point" would be the vertex

OpenStudy (sleepyhead314):

vertex form is: \(y=a(x-h)^2+k\) where \((h,k)\) is the vertex point and I forgot how to find a >,<

OpenStudy (iwanttogotostanford):

heres the pic of the graph

OpenStudy (iwanttogotostanford):

sorry can't get the pic

OpenStudy (iwanttogotostanford):

here are the answer choices though: y = (x − 4)2 − 4 y = (x + 4)2 − 4 y = (x + 2)2 + 6 y = (x + 2)2 + 12

OpenStudy (campbell_st):

well you know the vertex is (4, -4) so the equation is \[y = a(x -4)^2 - 4\] to find the value of a, substitute any point on the curve, other than the vertex, into the equation to find a

OpenStudy (iwanttogotostanford):

so it would be B?

OpenStudy (iwanttogotostanford):

help please guys @sleepyhead314 @wolf1728

OpenStudy (sleepyhead314):

take a look at what campbell_st wrote :)

OpenStudy (iwanttogotostanford):

i need the answer i still have to study for a test and i just need help please @sleepyhead314

OpenStudy (sleepyhead314):

campbell has showed clearly what your answer should be :)

OpenStudy (wolf1728):

campbell said the equation is y=a(x−4)^2−4

OpenStudy (iwanttogotostanford):

thank you guys, sorry a bit tired here